OpenStudy (anonymous):

HELP ME PLEASE!!! SECOND LAST QUESTION! The ratios ?:16 and 5:2 are equivalent. Find the missing number!!!!

OpenStudy (anonymous):

I will cry if don't get this answer! that's how much I despise numbers

OpenStudy (texaschic101):

ratios can be written as fractions ? : 16 is also x/16 5:2 is also 5/2 basically your dealing with equivalent fractions. 5/2....we want the denominators to be the same...so what is multiplied by 2 to get 16 ? answer is 8....what you multiply top number by, you have to multiply bottom number by the same number.. so...5/2 * 8/8 = 40/16 40/16 = 5/2 or 40:16 = 5:2